3 square root sign 64x7 divided by 3 square root sign 4x
Answer by checkley77(12844)

3sqrt64x^7/3sqrt4x factor the sqrt64=(8*8), the sqrtx^7=(x^3*x^3*x) & sqrt4=(2*2)
3*8x^3sqrtx/3*2sqrtx cancel out the sqrtx terms
24x^3/6 divide 24 by 6
4x^3 answer.